Fermi National Laboratory


Accelerator Update

Wednesday August 11
The day shift began with Tevatron (Tev) store #3723 colliding, stacking, and with beam to MiniBoone and Switchyard 120.

Shot Setup to the Recycler Ring (RR) began at 11:17 AM.

Shot setup is complete, we have returned to Antiproton source (Pbar) stacking at 12:30 PM.

At 7:00 PM the Antiproton Source (Pbar) stopped stacking and went to reverse protons for studies.

Pbar studies were completed at 11:34 PM, Pbar then returned to stacking.

Thursday August 12
The Midnight shift began with Store #3723 colliding, stacking, and with beam to MiniBoone, MCenter, and MTest.

A LCW water leak on the E3 bulk power supply resulted in the Tev store being terminated at 8:07 AM, two hours earlier then planned. Repairs then began on the E3 supply.

E3 bulk supply was repaired by 9:59 AM.

The Tev went into shot setup at 11:50 AM, with store #3725 colliding at 3:09 PM with an initial average luminosity of 91.71 E30.

Slipstacking studies began at 4:04 PM and were completed by 5:51 PM.

At 7:10 PM Pbar went into shot setup to the RR, this was completed by 7:55 PM with 34.5 mA injected into the RR.

At 8:05 PM the Antiproton Source (Pbar) went to reverse protons for studies.

Pbar studies were completed by 12:05 AM.

Friday August 13
The Midnight shift began with Store #3725 colliding, stacking, and with beam to MiniBoone and Switchyard 120.

Today's Plans
Continue to stack Pbars, some reverse proton studies and maintain the Tev store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 8/13/2004   email Fermilab